OpenStudy (anonymous):

find all parabolae which share the same tangent line at (1,1) as y=x^2

OpenStudy (anonymous):

maybe like this: general parabola formula is\[y=ax ^{2}+bx+c\] the general derivative is y'=2ax+b at x=1 y=x^2 derivative is y'=2 so parabolas for which 2a+b=2 will have same drivative

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw